K. Baggett, J. Benesch, M. Bickley, T. Carlino, G. Ciovati, J. Creel, E. Daly, K. Davis, M. Drury, C. Hovater, N. Huque, F. Humphrey, T. Ganey, C. Ginsburg, A. Kimber, L. King, G. Krafft, F. Marhauser, R. Michaud, C. Mounts, R. Nelson, W. Oren, A. Palczewski., T. Plawski, T. Powers, J. Preble, C. Reece, T. Reilly, A. Seryi, A. Solopova, M. Spata, and P. Vasilauskis
Agenda and Notes
1. Operations Status and Summary/Review (Bickley/Powers)
Run concluded successfully.
North Linac will be refurbished and most girders touched.
2. Progress on Particle Counter (Geng)
Particle Counter getting closer. Now design is focusing on "laser blanket" for aperture design. Detect phase shifts in interferometer arrangement.
July 22 updated solution and updated cost presented to division leadership. Waiting for favorable financial decision.
New Ion pump evaluation;